A body of mass m wants to leave completely from the earths gravitational attraction, find the minimum speed it requires at the earth surface (escape velocity)
1
Expert's answer
2021-01-31T15:11:46-0500
"KE=PE,""\dfrac{1}{2}mv^2=\dfrac{GM_Em}{R_E},""v_{escape}=\sqrt{\dfrac{2GM_E}{R_E}}."